OpenStudy (anonymous):

2x + 3y = 6 5x + 2y = 4 Which of the following equations could be the result of multiplication and addition to eliminate a variable in the system of equations? 19x = 24 19y = 22 11y = 22

OpenStudy (trojanpoem):

Multiply the first equation times -5 second times 2 -10x -15y = -30 10x + 4y = 8 -11y = -22 11 y= 22 (C) You can even solve for y and plug it in each. In more complex examples.

OpenStudy (anonymous):

thhank you:)

OpenStudy (trojanpoem):

Any time.